OpenStudy (anonymous):

A. Translate each of the following into algebraic expressions: (1 point each) 1.x plus y squared 2.8 times a number, x, decreased by y 3.the sum of x and y, squared 4.3 times the difference of x and y 5.6 less than x cubed B. Translate the following phrases into algebraic expressions. Explain, in complete sentences, the difference between the two. (3 points) •twice the sum of x and y •the sum of twice x and y C. Create your own example of a mathematical phrase you might hear in the real world and provide the translation for it. Be sure to identify what part of the phrase i

OpenStudy (anonymous):

A. 1. x+y^2 2.8ax-y 3.(x+y)^2 4.3(x-y) 5. 6<x^3 B. 1. 2(x+y) 2. 2x+y If we write the sum in parenthesis, it affects all thing in parenthesis. If we do not use parenthesis, the question is as it is.

OpenStudy (aroub):

Good answers! Just want to fix some answers. A. 2. 8x-y 5. x^3-6

OpenStudy (anonymous):

1. x plus y squared=x+y^2 2. 8 times a number, x, decreased by=8x-y 3. the sum of x and y, squared=(x+y)^2 4. 3 times the difference of x and y=3(x-y) 5. 6 less than x cubed=x^3-6 B. Translate the following phrases into algebraic expressions. Explain, in complete sentences, the difference between the two. (3 points) 1. twice the sum of x and y=2(x+y) 2. the sum of twice x and y=2x+y C. Create your own example of a mathematical phrase you might hear in the real world and provide the translation for it. Be sure to identify what part of the phrase is represented by your variable. (2 points) Im sorry but I don't know really get that.